The equation 2x2 + 3x + 1 = 0 has an irrational root. True/False. Please help me on this

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
2x2 + 3x + 1 = 0 ⇒ x = -1, -1/2 both are rational
∴ Statement is FALSE.
ALTERNATE SOLUTION :
If coefficients of a quadratic equation are rational and D is a perfect square, roots are rational.
As coefficients are rational and D = 9 – 8 = 1 is perfect square.
∴ Roots must be rational.
∴ Statement is FALSE.